OpenStudy (hannahp):

Help. Simplify, showing your steps. Tell the property you used in each step. 2(a + 4) + -8

OpenStudy (anonymous):

2a+8-8=0 2a=0 a=0/2 a=0

OpenStudy (hannahp):

Can you tell me the properties used in each step?

OpenStudy (anonymous):

This will help you out: http://www.math.com/school/subject2/lessons/S2U2L1DP.html